linda invested $12,000 for one year, part at 12 percent annual interest and the rest at 6 percent annual interest. how much did she invest at 6 percent to get $1,200 interest in a year?
Answer by FrankM(1040)

X is invested at 12%
12000-X is invested at 6%
.12X + .06(12000-X) = 1200
.12X + 720 - .06X = 1200
.06X = 480
X = $8000 ($960 interest from this portion)
12000-X = $4000 ($240 interest from this portion)